# On Emergent Social World Models — Evidence for Functional Integration of Theory of Mind and Pragmatic Reasoning in Language Models

### v1

@@ 3.2 Localization Procedure in LMs
For an autoregressive transformer model with a hidden dimension $d$, let $a_{l,t}\in\mathbb{R}^{d}$ be the output vector of each transformer block of the $l$-th layer at token $t$ of the input sequence (target or control stimulus). We record the outputs of each unit $i\in\{1,\dots,d\}$, which we call activations: $a_{l,t,i}$. To represent activations for the entire stimulus, we use activation at the last token before the model is required to produce the answer. Given a localizer dataset with a set of target stimuli $S_{T}$ and a set of control stimuli $S_{C}$, we generate two sets of activations for every unit $(l,i)$. We identify functional subnetworks based on relative differences in activation for $S_{T}$ in contrast to $S_{C}$, as measured by an unpaired or paired Welch’s $t$-test (the latter used for the simple suites GameBelief and MoralIntent). We write $\mathcal{T}(i,S_{T},S_{C})$ for the $t$-test statistic when comparing activations at unit $i$ for target stimuli in $S_{T}$ and control stimuli in $S_{C}$.
@@
We use two statistical approaches for localization. The simple approach defines a metric for unit $i$ under localizer $L$ as $m^{L}_{i}=\mathcal{T}\left(i,\bigcup_{j}S_{T}^{(j)},\bigcup_{k}S_{C}^{(k)}\right)$, i.e., the relevant $t$-test statistic for the union of all target and control items. We add to this a stricter conjunctive approach, which is common in cognitive neuroscience and uses so-called minimum statistic approach (Nichols et al., 2005) to detect significant differences between any pair of relevant target and control items. This metric for unit $i$ and suite $L$ selects the lowest $t$-statistic for all pairwise comparisons: $m^{L}_{i}=\min_{j,k}\mathcal{T}\left(i,S_{T}^{(j)},S_{C}^{(k)}\right)$. For simple suites, the conjunctive approach reduces to the simple.
@@
Based on localizer $L$, we select the subnetwork of units $i$, such that $m_{i}$ is statistically significant at the significance level $\alpha=0.05$ (modulo appropriate controlling for multiple comparisons). If that set comprises more than 1% of all units, we restrict the selection to the top 1% of units with the highest unique absolute measure $|m_{i}|$. For each localizer, we also construct a control condition based on the least active subnetwork by selecting the same number of units with the smallest absolute $|min_{i}|$-value from the subset of units whose $min_{i}$ is not significant. To investigate the causal role of the subnetworks through ablation experiments (Section 3.5), units in the subnetwork are set to zero.
@@
Altogether, we consider eight localizers for identifying relevant subnetworks for subsequent analyses. We have five simple localizers, four of which are obtained from each of the the four localizer suites, and another one from the union of all control and target sets from all localizer suites (the all localizer). We consider three conjunctive localizers, two of which come from the LatentBeliefs and the CommunicativeIntent suites. A third conjunctive localizer was constructed from the union of these two suites (the LB + CI localizer), as they have non-paired stimuli and were considered to be most relevant for pragmatic reasoning.
@@ 3.5 Ablating Localized Subnetworks
To assess the causal role of the localized networks, we conducted ablation experiments with all eight localizers, evaluating 20 models on selected datasets (see Section 3.3) with as in Section 2.2.
We test predictions about three causal effects (on ToM, on pragmatics and on syntax), which, if observed, conjunctively provide evidence for functional integration (see the upper panel of Figure 3):
@@
    ToM-network ablation decreases performance on ToM-related tasks
@@
    causal effect (CE) on ToM-task performance is bigger for ToM ablation than for control
@@
    ToM-network ablation decreases performance on pragmatic tasks
@@
    CE on pragmatic abilities is bigger for ToM ablation than for control
@@
    ToM-network ablation does not decrease performance on syntax (BLiMP)
@@
    CE is bigger on pragmatics than on syntax
@@
These predictions can be tested at different levels of aggregation. For example, results for some localizers may conform to our predictions better than results from others (see the two right-most plots on the bottom row of Figure 3). In light of such variation, the most insightful conclusions are those from a “global analysis,” which marginalizes over all LMs and all localization methods and whose results are shown in the left-most facet in the bottom row of Figure 3. In order to test predictions from the “global analysis” we inspect the relevant (marginalized) posterior probabilities from a Bayesian beta regression model. [^6] Model in R syntax: accuracy $\sim$ model_type + model_size + ablation_localizer * domain, where ablation_localizer is a variable with 17 levels combining the two ablations and eight localizations procedures and the baseline intact condition. Marginalizing across localizers and models, both P1.1 ($\beta=0.24[0.14,0.35]$) and P1.2 ($\beta=0.07[0.02,0.11]$) are credibly borne out in our results, providing evidence for a general causal role of the localized ToM networks for LMs’ ToM performance. Similarly, both P2.1 ($\beta=0.30[0.21,0.40]$) and P2.2 ($\beta=0.07[0.03,0.11]$) were credibly borne out, indicating that the ToM subnetworks were generally causally involved in LMs’ pragmatic performance. Finally, the data supports P3.1 (no credible decrease of syntax performance under ToM ablation: $\beta=0.15[-0.17,0.49]$), but not P3.2 ($\beta=0.15[-0.20,0.48]$), offering mild evidence that the causal role of the ToM subnetworks is stronger for pragmatic than syntactic capabilities, from the “global” point of view. Taken together, our results lend credence to the idea that ToM and pragmatic reasoning do share causally relevant functional units to a non-trivial degree.
@@
Towards further exploration, when testing predictions separately for each of the eight localizers, we observe that, across models, predictions 1 and 2 were borne out only for LatentBeliefs (simple & conjunctive) and GameBeliefs (simple), suggesting that these drive the main results above (shown in Figure 10, full statistical results presented in Table 5 in the Appendix). For all other localizers (except for CommunicativeIntent (conjunctive)), only P1.1, P2.1 and P3.1 were borne out. We note that the three localizers supporting our predictions in full also showed similar distributional patterns in the models (Figure 2), different from other localizers. Results for all localizers and all models are shown in the Appendix in Figures 1213.
@@
For better understanding of the localization methods, we compare the causal effects of ablations of subnetworks localized with the simple vs. the conjunctive method for LatentBeliefs (LB) and CommunicativeIntent (CI). Aggregating causal effects on ToM and pragmatics relative to the baseline (P1.1 and P2.1), we find no credible differences for LB ($\beta=0.06[-0.03,0.15]$), nor CI ($\beta=0.06[-0.03,0.15]$). Relative to the control ablation (P1.2 and P2.2), we also find no credible differences (LB: $\beta=0.10[-0.03,0.23]$, CI: $\beta=0.07,[-0.06,0.20]$). The proportion of the units localized by both methods is moderately high (LB = 0.50, CI = 0.66). This suggests that (i) simple and conjunctive strategies may be interchangeable, and (ii) the obserserved causal effects may be due to only a subset of the localized subnetworks.
@@
Additionally, we compare the causal effects of the ToM ablation between different model sizes using a beta regression model with a three-way interaction. [^7] Model in R syntax: accuracy $\sim$ model_type + model_size * ablation_localizer * domain. Averaged across localizers, we found no credible differences (L vs. M models: ToM: $\beta=0.15[-0.12,0.42]$, pragmatics: $\beta=0.13[-0.13,0.39]$; M vs. S models: ToM: $\beta=-0.04[-0.27,0.19]$, pragmatics: $\beta=-0.04[-0.46,0.39]$).
@@ 4 Discussion
This paper offered five contributions towards addressing the overarching research question of whether LMs contain “social world models” — latent representations of information related to social reasoning about other minds. (1) We outlined the functional integration hypothesis, according to which internal representations and computational mechanisms are shared between ToM and pragmatic reasoning tasks, and we derived concrete, testable predictions from this view. (2) We report stringent theory-driven statistical analyses based on data from both behavioral evaluations and causal ablation studies which provide suggestive evidence in favor of the functional integration view. (3) Improving on previous work (AlKhamissi et al., 2025a), this paper contributes four novel synthetically generated localization suites with a total of 1400 stimuli based on localizers carefully assembled from prior work in cognitive neuroscience that spans a substantial swath of various aspects of ToM reasoning. (4) We ported established methods from neuroscience (the conjunctive analysis) into the realm of language models, and presented first results to assess their value. (5) Finally, we explored a finer-grained analysis of ToM based on the ATOMS framework and found that its subcategories are indeed predictive of benchmark accuracies. They also help to interpret the results from the causal ablation studies, where the ablation of localized subnetworks addressing those ATOMS categories that were part of the best explanations of the behavioral data had the strongest causal effect of interest on LMs’ ToM and pragmatic capabilities, although it remains puzzling why ablating the subnetwork identified by the CommunicativeIntent localizer did not have the predicted causal effect.
@@
Even though we found support for various causal effects predicted by a functional integration view, our results hinge on whether the datasets used for localizations and behavioral analysis indeed operationalize the right kinds of concepts (general ToM, specific aspects in the ATOMS framework, pragmatic reasoning etc.). As is the case with human experiments, the validity of methods and operations require converging evidence from many like-minded investigations and replications. What is more, while it is commonly assumed that ToM and pragmatics tasks recruit reasoning about mental causes, we should be open to the possibility that, also in humans, explicit social reasoning trades off with cheaper heuristics Rubio-Fernandez (2024). Together, these considerations underline the importance of an interdisciplinary perspective on language abilities, closely comparing human and machine solutions to superficially similar tasks. In this spirit, the present paper contributes by investigating LMs’ functional mechanisms of different aspects of social reasoning at a high resolution, offering exciting avenues for future investigations of the exact relation between different capabilities and of the architectural pressures shaping LMs’ “social world models.”
@@
The presented functional localization and evaluations make a number of design choices in the procedure which are known to potentially impact the results (Bommasani et al., 2023), and future work should evaluate the robustness of the results to such design choices.
For the baseline and ablation behavioral evaluations, we choose to use means over the conditional log probabilities of the answer tokens. For the subnetwork localization, we used one general instruction and one specific prompt format across stimuli. All localizer stimuli contained a forced-choice task, which was formulated as a sentence continuation task with two answer options for all localizer suites, except for CommunicativeIntent, for which the task was a meta-linguistic judgment task to select whether an utterance was literal, ironic, deceitful or meaningless.
@@
Furthermore, future work could expand the selection of datasets used to evaluate the syntactic abilities of models beyond the BLiMP dataset. Additionally, all datasets focus exclusively on English, a resource-rich language; in particular, pragmatics datasets in other languages may differ with respect to the cultural and social norms that must be reasoned about. The ToM datasets contain a large proportion of false-belief tasks, whereas real-world language use may provide more evidence for true-belief-based reasoning, which may therefore be strongly represented in LMs’ training data. Finally, since some of our datasets have been publicly available for several years, there is a risk that the training data of recent models are contaminated with these datasets.
Although the synthetic generation of novel localizer materials was based on validated materials from neuroscience research and allowed for a larger-scale study with language models, these new materials have not been validated in human fMRI studies. Therefore, while drawing inspiration from cognitive neuroscience research, our conclusions should not be taken as claims about human-likeness of the functional organization of LMs.
@@ Acknowledgments
JFK, AM anf MF were supported by the Volkswagen Foundation through a Momentum grant. MF is a member of the Machine Learning Cluster of Excellence at University of Tübingen, EXC number 2064/2 – Project number 39072764. PT and MF gratefully acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG, and the data storage service SDS@hd supported by the Ministry of Science, Research and the Arts Baden-Württemberg (MWK) and the German Research Foundation (DFG) through grant INST 35/1803-1 FUGG and INST 35/1804-1 LAGG.

### v2

@@ 3.2 Localization Procedure in LMs
For an autoregressive transformer model with a hidden dimension $d$, let $a_{l,t}\in\mathbb{R}^{d}$ be the output vector of each transformer block of the $l$-th layer at token $t$ of the input sequence (target or control stimulus). We record the outputs of each unit $i\in\{1,\dots,d\}$, which we call activations: $a_{l,t,i}$. To represent activations for the entire stimulus, we use activation at the last token before the model is required to produce the answer. Given a localizer dataset with a set of target stimuli $S_{T}$ and a set of control stimuli $S_{C}$, we generate two sets of activations for every unit $(l,i)$ (i.e., target and control activations). We identify functional subnetworks based on relative differences in activation for $S_{T}$ in contrast to $S_{C}$, as measured by an unpaired or paired Welch’s $t$-test (the latter used for the simple suites GameBelief and MoralIntent). We write $\mathcal{T}(i,S_{T},S_{C})$ for the $t$-test statistic when comparing activations at unit $i$ for target stimuli in $S_{T}$ and control stimuli in $S_{C}$.
@@
We use two statistical approaches for localization. The simple approach defines a metric for unit $i$ under localizer $L$ as $m^{L}_{i}=\mathcal{T}\left(i,\bigcup_{j}S_{T}^{(j)},\bigcup_{k}S_{C}^{(k)}\right)$, i.e., the relevant $t$-test statistic for the union of all target and control items. We add a stricter conjunctive approach, which is common in cognitive neuroscience and uses a so-called minimum statistic approach (Nichols et al., 2005) to detect significant differences between any pair of relevant target and control items. This metric for unit $i$ and suite $L$ selects the lowest $t$-statistic for all pairwise comparisons: $m^{L}_{i}=\min_{j,k}\mathcal{T}\left(i,S_{T}^{(j)},S_{C}^{(k)}\right)$. For simple suites, the conjunctive approach reduces to the simple.
@@
Based on localizer $L$, we select the subnetwork of units $i$, such that $m_{i}$ is statistically significant at the significance level $\alpha=0.05$ (modulo appropriate controlling for multiple comparisons). If that set comprises more than 1% of all units, we restrict the selection to the top 1% of units with the highest unique absolute measure $|m_{i}^{L}|$. For critical ToM localization with each suite $L$, we also construct a control localization condition based on the least active subnetwork by selecting the same number of units with the smallest absolute $|m_{i}^{L}|$-value from the subset of units whose $m_{i}^{L}$ is not significant. To investigate the causal role of the localized subnetworks through ablation experiments (Section 3.5), units in the subnetwork are set to zero.
@@
Altogether, we consider eight localizers for identifying relevant subnetworks for subsequent analyses. We have five simple localizers, four of which are obtained from each of the the four localizer suites, and another one from the union of all target and control sets from all localizer suites (the all localizer). We consider three conjunctive localizers, two of which come from the LatentBeliefs and the CommunicativeIntent suites. A third conjunctive localizer was constructed from the union of these two suites (the LB + CI localizer), as they have non-paired stimuli and were considered to be most relevant for pragmatic reasoning.
@@ 3.5 Ablating Localized Subnetworks
To assess the causal role of the localized networks, we first ablated, for all 20 models and each localizer $L$, the ToM network identified by $L$ and the “least active” control network for $L$, and then evaluated the performance on selected datasets as described in Section 2.2 for each ablation.
We test predictions about three causal effects (on ToM, on pragmatics and on baseline general linguistic abilities (BLiMP and SNLI)), which, if observed, conjunctively provide evidence for functional integration (see the upper panel of Figure 3):
@@
    ToM-network ablation decreases performance on ToM-related tasks.
@@
    CE (causal effect) on ToM-task performance is bigger for ToM ablation than for control.
@@
    ToM-network ablation decreases performance on pragmatic tasks.
@@
    CE on pragmatic abilities is bigger for ToM ablation than for control ablation.
@@
    ToM-network ablation does not decrease performance on baseline linguistic benchmarks.
@@
    CE is bigger on pragmatics than on baseline.
@@
These predictions can be tested at different levels of aggregation. For example, results for some localizers may conform to our predictions better than results from others (see the two right-most plots in the bottom row of Figure 3). In light of such variation, the most insightful conclusions are those from a “global analysis,” which marginalizes over all LMs and all localization methods and whose results are shown in the left-most facet in the bottom row of Figure 3. In order to test predictions from the “global analysis” we inspect the relevant posterior probabilities from a Bayesian beta regression model. [^6] Model in R syntax: accuracy $\sim$ model_type + model_size + ablation_localizer * domain, where ablation_localizer is a variable with 17 levels combining the two ablations and eight localizations procedures and the baseline intact condition (results from Section 2.3). Marginalizing across localizers and models, both P1.1 ($\beta=0.25[0.14,0.35]$) and P1.2 ($\beta=0.06[0.02,0.11]$) were credibly borne out in our results, providing evidence for a general causal role of the localized ToM networks for LMs’ ToM performance. Similarly, both P2.1 ($\beta=0.30[0.20,0.39]$) and P2.2 ($\beta=0.07[0.02,0.12]$) were credibly borne out, indicating that the ToM subnetworks were generally causally involved in LMs’ pragmatic performance. Finally, the data supported P3.1 (no credible decrease of linguistic benchmark performance under ToM ablation: $\beta=0.13[-0.10,0.35]$), but not P3.2 ($\beta=0.17[-0.07,0.41]$), offering mild evidence that the causal role of the ToM subnetworks is stronger for pragmatic than general linguistic capabilities, from the “global” point of view. Together, our results support to the idea that ToM and pragmatic reasoning do share causally relevant functional units in LMs to a non-trivial degree.
@@
Towards further exploration, we tested predictions separately for each of the eight localizers. Across models, predictions 1.1, 2.1, and 3.1 were borne out for all localizers (except for CommunicativeIntent (conjunctive) (1.1) and GameBeliefs (3.1)). Predictions 1.2 and 2.2 were borne out only for LatentBeliefs (simple & marginally conjunctive) and GameBeliefs, suggesting that these three localizers mainly drive the results above (shown in Figure 14). This suggests that subnetworks tapping into the ToM aspects beliefs, percepts, desires and emotions (captured by these three localizers) might functionally support both ToM and pragmatic performance of LMs. We note that these three localizers also showed similar distributional patterns in the models (Figure 2), different from other localizers. Results for all localizers and models are shown in the Appendix in Figures 1617, full statistical results are presented in Table 5.
Moreover, we investigated to which extent the identified subnetworks functionally support mental state related reasoning as opposed to other cognitive skills that may be recruited in ToM and pragmatic tasks. In addition to the general linguistic ability benchmarks, we evaluated selected models on the entity tracking (Kim and Schuster, 2023) and analogical reasoning tasks (Webb et al., 2023). When ablating the critical and control subnetworks, e.g., localized by LatentBeliefs and GameBeliefs, we found that the critical ablations had a similar effect on entity tracking performance as on ToM and pragmatic performance across models, but did not as consistently affect other datasets, suggesting that ToM and pragmatic capabilities might indeed also rely on shared entity tracking mechanisms. Details of the analysis are reported in Appendix D.1.
@@
For better understanding the impact of the statistical methods for localization, we compared the causal effects of ablations of subnetworks localized with the simple vs. the conjunctive method for LatentBeliefs (LB) and CommunicativeIntent (CI). Aggregating causal effects on ToM and pragmatics relative to the baseline (P1.1 and P2.1), we found no credible differences for LB ($\beta=0.05[-0.03,0.15]$), nor CI ($\beta=0.06,[-0.03,0.15]$). Relative to the control ablation (P1.2 and P2.2), we also found no credible differences (LB: $\beta=0.10[-0.03,0.23]$, CI: $\beta=0.07,[-0.06,0.20]$). The proportion of the units localized by both methods is moderately high (LB = 0.50, CI = 0.66). This suggests that (i) simple and conjunctive strategies may be interchangeable, and (ii) the observed causal effects may be due to only a subset of the localized subnetworks.
@@
Finally, we compared the causal effects of the ToM ablation between different model sizes using a beta regression model with a three-way interaction. [^7] Model in R syntax: accuracy $\sim$ model_type + model_size * ablation_localizer * domain. Averaged across localizers, we found no credible differences in the causal effect on ToM and pragmatic performance of models of different sizes (L vs. M models: ToM: $\beta=0.15[-0.12,0.42]$, pragmatics: $\beta=0.13[-0.12,0.38]$; M vs. S models: ToM: $\beta=-0.04[-0.26,0.18]$, pragmatics: $\beta=-0.11[-0.41,0.20]$), suggesting that varying capacity of the models may not directly affect functional integration.
@@ 4 Discussion
This paper offered five contributions towards addressing the overarching research question of whether LMs contain “social world models” — latent representations of information related to social reasoning about other minds. (1) We outlined the functional integration hypothesis, according to which internal representations and computational mechanisms are shared between ToM and pragmatic reasoning tasks, and we derived concrete, testable predictions from this view. (2) We report stringent theory-driven statistical analyses based on data from both behavioral evaluations and causal ablation studies which provide suggestive evidence in favor of the functional integration view. (3) Improving on previous work (AlKhamissi et al., 2025a), this paper contributes four novel synthetically generated localization suites with a total of 1400 stimuli based on localizers carefully assembled from prior work in cognitive neuroscience that spans a substantial swath of various aspects of ToM reasoning. (4) We ported established methods from neuroscience (the conjunctive analysis) into the realm of language models, and presented first results to assess their value. (5) Finally, we explored a finer-grained analysis of ToM based on the ATOMS framework and found that its subcategories are indeed predictive of benchmark accuracies. They also help to interpret the results from the causal ablation studies, where the ablation of localized subnetworks addressing those ATOMS categories that were part of the best explanations of the behavioral data also had the strongest causal effect of interest on LMs’ ToM and pragmatic capabilities, although it remains puzzling why ablating the subnetwork identified by the CommunicativeIntent localizer did not have the predicted causal effect.
@@
Even though we found support for various causal effects predicted by a functional integration view, our results hinge on whether the datasets used for localizations and behavioral analysis indeed operationalize the right kinds of concepts (general ToM, specific aspects in the ATOMS framework, pragmatic reasoning etc.). As is the case with human experiments, the validity of methods and operations require converging evidence from many like-minded investigations and replications. What is more, while it is commonly assumed that ToM and pragmatics tasks recruit reasoning about mental causes, we should be open to the possibility that, also in humans, explicit social reasoning trades off with cheaper heuristics Rubio-Fernandez (2024). Additionally, both tasks may share not only social reasoning, but likely also other general intelligence capabilities. While the exact nature of such general mechanisms involved in ToM and pragmatics remains elusive (for both humans and machines), our novel large localizer datasets and fine-grained statistical methodology for establishing functional integration between such tasks show one potential path towards contributing to this deep issue.
Together, these considerations underline the importance of an interdisciplinary perspective on language abilities, closely comparing human and machine solutions to superficially similar tasks. In this spirit, the present paper contributes by investigating LMs’ functional mechanisms of different aspects of social reasoning at a high resolution, offering exciting avenues for future investigations of the exact relation between different capabilities and of the architectural pressures shaping LMs’ “social world models.”
@@
The presented functional localizations and evaluations make a number of design choices in the procedure which are known to potentially impact the results (Bommasani et al., 2023), and future work should evaluate the robustness of the results to such design choices.
For the baseline and ablation behavioral evaluations, we chose to use means over the conditional log probabilities of the answer tokens. For the subnetwork localization, we used one general instruction and one specific prompt format across stimuli. All localizer stimuli contained a forced-choice task, which was formulated as a sentence continuation task with two answer options for all localizer suites, except for CommunicativeIntent, for which the task was a meta-linguistic judgment task to select whether an utterance was literal, ironic, deceitful or meaningless.
@@
Furthermore, future work could expand the selection of datasets used to evaluate the general linguistic abilities of models beyond the BLiMP and SNLI datasets. Additionally, all datasets focus exclusively on English, a resource-rich language; in particular, pragmatics datasets in other languages may differ with respect to the cultural and social norms that must be reasoned about. The precision and diversity of the pragmatics datasets may also vary between the selected benchmarks, e.g., due to semi-automated construction of some of the materials. The ToM datasets contain a large proportion of false-belief tasks, whereas real-world language use may provide more evidence for true-belief-based reasoning, which may therefore be strongly represented in LMs’ training data. Finally, since some of our datasets have been publicly available for several years, there is a risk that the training data of recent models are contaminated with these datasets.
Although the synthetic generation of novel localizer materials was based on validated materials from neuroscience research and allowed studying language models at a larger scale than related previous work, these new materials have not been validated in human fMRI studies. Therefore, while drawing inspiration from cognitive neuroscience research, our conclusions cannot be directly translated to claims about human-likeness of the functional organization of LMs.
@@ Acknowledgments
JFK, AM and MF were supported by the Volkswagen Foundation through a Momentum grant. MF is a member of the Machine Learning Cluster of Excellence at University of Tübingen, EXC number 2064/2 – Project number 39072764. PT and MF gratefully acknowledge support by the state of Baden- Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG, and the data storage service SDS@hd supported by the Ministry of Science, Research and the Arts Baden-Württemberg (MWK) and the German Research Foundation (DFG) through grant INST 35/1803-1 FUGG and INST 35/1804-1 LAGG.